# Know When to Abstain: Optimal Selective Classification with Likelihood Ratios

### v1

@@ 5 Experiments
Datasets.
We evaluate our methods across vision and language domains, with a primary focus on the former. For vision tasks, we use ImageNet-1K (Im-1K) and a suite of covariate-shifted variants: 1) ImageNet-Rendition (Im-R) [31], 2) ImageNet-A (Im-A) [32], 3) ObjectNet (ON) [33], 4) ImageNetV2 (Im-V2) [34], 5) ImageNet-Sketch (Im-S) [35], and 6) ImageNet-C (Im-C) [36]. We group these datasets based on label coverage: full 1000-class coverage (Im-1K, Im-V2, Im-S, Im-C) and subsets of classes (Im-R, Im-A, ON). For language tasks, we evaluate on the Amazon Reviews dataset [37, 38]. To simulate realistic deployment scenarios involving distribution shift, following Liang et al. [3] we evaluate on mixed test sets that combine in-distribution and covariate-shifted samples. For example, results reported on Im-C are computed on a combined test set of Im-1K and Im-C.
Classifiers and Baseline Selector Scores.
For vision experiments, we consider two families of classifiers, CLIP-based zero-shot VLMs [39] and supervised classifiers. Specifically, we use the CLIP model from Data Filtering Networks (DFN) [40] and EVA [41] for supervised learning, chosen for their state-of-the-art accuracy on ImageNet. Our focus is not on model training but on evaluating selector scores applied post-hoc. Note that for EVA, we restrict evaluation to datasets with full 1K class coverage as the model is trained on the complete ImageNet label set only. In contrast, CLIP can be adapted at inference time to arbitrary label subsets, so we evaluate it across all datasets. For language tasks, we fine-tune a DistilBERT [42] model using LISA [43] on the Amazon Reviews training set and evaluate selective classification performance on the full test set.
In terms of baseline selector scores, we compare our proposed selector scores $\Delta$-MDS, $\Delta$-KNN, and their linear combinations with commonly used OOD detection and uncertainty-based scores: MSP [2], MCM (for CLIP), MaxLogit [24], Energy [25], MDS [5], KNN [6], and RLog [3]. As they are functionally similar, we will abbreviate MCM as MSP when presenting CLIP results for simplicity. Detailed experimental settings are provided in Appendix D.
Evaluation Metrics.
We evaluate performance using two metrics: the Area Under the Risk-Coverage Curve (AURC) and the Normalized AURC (NAURC) [17]. The AURC captures the joint performance of the classifier and selector across coverage levels. NAURC normalizes AURC to account for the classifier’s base error rate, providing a fairer comparison across models with different accuracies. Formally, NAURC is defined as:
where $g^{*}$ denotes an oracle confidence function achieving optimal AURC, and $R(f)$ is the risk of $f$. Intuitively, NAURC measures how close the selector $g$ gets to the optimal, normalized by the classifier’s total error. Thus, while AURC is useful for understanding overall performance in the context of a specific model, NAURC enables fair selector comparisons across models by factoring out baseline classifier accuracy.
5.1 Image Experiments
@@
Figure 2: Risk-coverage curves of various selector methods for CLIP (top row) and EVA (bottom row).
@@
We report full selective classification results for CLIP and EVA models in Table 1 and Table 2, respectively. First, let us consider CLIP results. We see that going from MDS and KNN to their NP-informed variants, $\Delta$-MDS and $\Delta$-KNN, leads to roughly $50\%$ reduction in average AURC and NAURC, showing that the assumptions made in the NP-optimality theory hold well in practice. The best average performance is achieved by the linear combinations $\Delta$-KNN-RLog and $\Delta$-MDS-RLog, with $\Delta$-KNN-RLog leading overall in both AURC and NAURC. RLog score ranks third on average, highlighting its strength as a standalone logit-based selector. Motivated by this strong performance, we use RLog in combination with our distance-based scores. Figure 2 (top row) illustrates the risk-coverage curves for selected datasets for CLIP. $\Delta$-KNN-RLog consistently demonstrates the most favorable trade-off across all coverage levels, remaining stable even at low coverage. In contrast, methods such as MDS and Energy suffer from sharp risk increases at low coverage, likely due to the incorrect rejection of samples that contribute disproportionately to risk at low coverage.
For EVA, the ranking is reversed: $\Delta$-MDS-RLog achieves the best overall performance, followed by $\Delta$-KNN-RLog. This supports our hypothesis that MDS-based methods are particularly effective for supervised models due to the close connection between softmax classifiers and Gaussian Discriminant Analysis [5], which justifies the Gaussian assumptions used in MDS. In contrast, CLIP models trained with contrastive learning [39] do not satisfy these assumptions, making the nonparametric $\Delta$-KNN combination more suitable. Figure 2 (bottom row) confirms that $\Delta$-MDS-RLog yields the best risk-coverage behavior for supervised learning across all coverage levels. Comparing average NAURC on the full 1K-class datasets, $\Delta$-MDS-RLog with EVA is the top performing selector score, with a slightly better score (0.157) than the best performer on CLIP (0.163).
@@
For practitioners aiming to identify the best overall selective classification setup that considers both the base classifier and the selector, a practical approach is to compare performance using the AURC metric. On the 1K-class datasets, EVA paired with $\Delta$-MDS-RLog achieves an AURC of 3.86, outperforming the DFN CLIP model with $\Delta$-KNN-RLog at 5.01. Despite similar NAURC values, EVA’s higher Im-1K base accuracy ($84.33\%$ vs. DFN CLIP’s $80.39\%$) makes it the preferred choice when considering both components. Intuitively, the optimal setup involves pairing the best selector (here, $\Delta$-MDS-RLog) with the most accurate base classifier.
@@ 5.2 Language Experiments
Table 3: Results on Amazon Reviews dataset and its covariate shifted test set at full coverage using DistilBERT trained with LISA.
In-D Cov Shift
Method A N A N
MSP 12.2 0.368 13.9 0.401
MaxLogit 12.6 0.384 14.3 0.416
Energy 12.89 0.397 14.6 0.428
MDS 20.6 0.739 22.2 0.750
KNN 19.4 0.686 21.3 0.711
RLog 12.4 0.376 14.1 0.410
$\Delta$-MDS 12.7 0.389 14.4 0.422
$\Delta$-KNN 12.4 0.374 14.2 0.412
$\Delta$-MDS-RLog 12.2 0.368 13.9 0.401
$\Delta$-KNN-RLog 12.0 0.358 13.8 0.394
$\Delta$-MDS-MSP 11.9 0.354 13.6 0.387
$\Delta$-KNN-MSP 12.0 0.359 13.8 0.396
Table 3 presents results on the Amazon Reviews dataset. Unlike the vision tasks, the best-performing method is $\Delta$-MDS-MSP, followed closely by $\Delta$-MDS-RLog and $\Delta$-KNN-MSP. Since LISA [43] uses a softmax classification objective, the superiority of MDS-based selectors supports our hypothesis about their suitability for supervised models. Interestingly, MSP outperforms RLog in this domain, resulting in better performance when combined with $\Delta$-MDS. This highlights another important practical insight that the best linear combination often involves pairing the top-performing standalone distance-based and logit-based score.
@@ 5.3 Ablations
Table 4 summarizes several ablation experiments. First, we justify averaging the top-$k$ nearest neighbor distances in $\Delta$-KNN rather than using the $k$-th distance alone. This modification yields measurable gains, where average AURC improves from 6.03 to 5.89 and NAURC improves from 0.233 to 0.225. We also investigate various combinations of selector scores. For CLIP, $\Delta$-KNN-RLog remains the best across all configurations, outperforming both double-distance combinations (e.g., $\Delta$-MDS-$\Delta$-KNN) and double-logit combinations (e.g., MSP-RLog). Notably, pairing $\Delta$-KNN with RLog significantly outperforms pairing it with MSP, further validating RLog’s role as a strong logit-based complement.
@@ 6 Conclusion
We presented a principled framework for designing selector functions for selective classification, grounded in the Neyman–Pearson lemma. This perspective reveals that the optimal selection score is a monotonic transformation of a likelihood ratio, which is a unifying insight that explains the behavior of several existing methods. Building on this foundation, we proposed two novel distance-based scores and their linear combinations with logit-based baselines. Extensive experiments across vision and language tasks under covariate shift demonstrate that our methods achieve state-of-the-art performance across diverse settings.
@@
While our focus has been on classification, the Neyman–Pearson framework is general and broadly applicable to other predictive tasks. Exploring selective prediction in settings where uncertainty plays a critical role, such as regression, semantic segmentation, or time series forecasting, presents promising future directions. Additionally, extending these ideas to generative models is especially timely. Large language models (LLMs), for instance, are prone to hallucinations [44], where the model confidently generates incorrect information. Enabling LLMs to recognize uncertainty and selectively abstain from answering, in other words selective generation, offers another exciting avenue for future work.
7 Acknowledgements
This research is supported by A*STAR under its National Robotics Programme (NRP) (Award M23NBK0053).
@@ 7 Acknowledgements
Appendix A Proofs

### v2

@@ 5 Experiments
Datasets.
We evaluate our methods across vision and language domains, with a primary focus on the former. For vision tasks, we use ImageNet-1K (Im-1K) and a suite of covariate-shifted variants: 1) ImageNet-Rendition (Im-R) (Hendrycks et al., 2020), 2) ImageNet-A (Im-A) (Hendrycks et al., 2021), 3) ObjectNet (ON) (Barbu et al., 2019), 4) ImageNetV2 (Im-V2) (Recht et al., 2019), 5) ImageNet-Sketch (Im-S) (Wang et al., 2019), and 6) ImageNet-C (Im-C) (Hendrycks and Dietterich, 2019). We group these datasets based on label coverage: full 1000-class coverage (Im-1K, Im-V2, Im-S, Im-C) and subsets of classes (Im-R, Im-A, ON). For language tasks, we evaluate on the Amazon Reviews dataset (Ni et al., 2019; Koh et al., 2021). To simulate realistic deployment scenarios involving distribution shift, following Liang et al. (2024) we evaluate on mixed test sets that combine in-distribution and covariate-shifted samples. For example, results reported on Im-C are computed on a combined test set of Im-1K and Im-C.
Classifiers and Baseline Selector Scores.
We consider two families of classifiers for vision experiments, namely CLIP zero-shot VLMs (Radford et al., 2021) and supervised classifiers. Specifically, we use the CLIP model from Data Filtering Networks (DFN) (Fang et al., 2024) and EVA (Fang et al., 2023) for supervised learning, chosen for their state-of-the-art accuracy on ImageNet. Our focus is not on model training but on evaluating selector scores applied post-hoc. Note that for EVA, we restrict evaluation to datasets with full 1K class coverage as the model is trained on the complete ImageNet label set only. In contrast, CLIP can be adapted at inference time to arbitrary label subsets, so we evaluate it across all datasets. For language tasks, we fine-tune a DistilBERT (Sanh et al., 2019) model using LISA (Yao et al., 2022) on the Amazon Reviews training set and evaluate selective classification performance on the full test set.
For baseline scores, we compare our proposed $\Delta$-MDS, $\Delta$-KNN, and their linear combinations with common OOD detection and uncertainty-based scores: MSP (Hendrycks and Gimpel, 2017), MCM (for CLIP), MaxLogit (Hendrycks et al., 2019), Energy (Liu et al., 2020), MDS (Lee et al., 2018), KNN (Sun et al., 2022), and RLog (Liang et al., 2024) as well as SIRC (Xia and Bouganis, 2022). As they are functionally similar, we abbreviate MCM as MSP when presenting CLIP results. Details of the baseline are provided in Appendix A.
Evaluation Metrics.
We evaluate performance using two metrics: the Area Under the Risk-Coverage Curve (AURC) and the Normalized AURC (NAURC) (Cattelan and Silva, 2023). The AURC captures the joint performance of the classifier and selector across coverage levels. NAURC normalizes AURC to account for the classifier’s base error rate, providing a fairer comparison across models with different accuracies. Formally, NAURC is defined as:
where $g^{*}$ denotes an oracle confidence function achieving optimal AURC, and $R(f)$ is the risk of $f$. The oracle can be computed in practice using the ground-truth labels of the evaluation set. Intuitively, NAURC measures how close the selector $g$ gets to the optimal, normalized by the classifier’s total error. Thus, while AURC is useful for understanding overall performance in the context of a specific model, NAURC enables fair selector comparisons across models by factoring out baseline classifier accuracy.
Table 2: Results on Amazon Reviews and its covariate shifted test set at full coverage using DistilBERT trained with LISA.
In-D Cov Shift
Method A N A N
MSP 12.2 0.368 13.9 0.401
MaxLogit 12.6 0.384 14.3 0.416
Energy 12.89 0.397 14.6 0.428
MDS 20.6 0.739 22.2 0.750
KNN 19.4 0.686 21.3 0.711
RLog 12.4 0.376 14.1 0.410
SIRC 12.3 0.370 14.0 0.403
$\Delta$-MDS 12.7 0.389 14.4 0.422
$\Delta$-KNN 12.4 0.374 14.2 0.412
$\Delta$-MDS-RLog 12.2 0.368 13.9 0.401
$\Delta$-KNN-RLog 12.0 0.358 13.8 0.394
$\Delta$-MDS-MSP 11.9 0.354 13.6 0.387
$\Delta$-KNN-MSP 12.0 0.359 13.8 0.396
Selecting $\lambda$ and $k$.
Both $\lambda$ and $k$ can be selected on a validation set. We found that the simplest recipe to fitting $\lambda$ is to balance the magnitudes of $s_{1}({\bm{x}})$ and $s_{2}({\bm{x}})$, so that neither overpowers the other. For $k$ in KNN-based scores, we find that $k\in[25,50]$ is a sweet spot. Full experimental settings are provided in Appendix E, and hyperparameter sensitivity analysis for $\lambda$ and $k$ is presented in appendix Fig. 2.
5.1 Image Experiments
@@
We report full selective classification results for CLIP and EVA models in Table 1 and Table 3, respectively. First, let us consider CLIP results. We see that going from MDS and KNN to their NP-informed variants, $\Delta$-MDS and $\Delta$-KNN, leads to roughly $50\%$ reduction in average AURC and NAURC, showing that the assumptions made in the NP-optimality theory hold well in practice. The best average performance is achieved by the linear combinations $\Delta$-KNN-RLog and $\Delta$-MDS-RLog, with $\Delta$-KNN-RLog leading overall in both AURC and NAURC. RLog score ranks third on average, highlighting its strength as a standalone logit-based selector. Motivated by this strong performance, we use RLog in combination with our distance-based scores. We plot the risk-coverage curves for selected datasets in Fig. 3 of the appendix, showing our methods consistently demonstrates the most favorable trade-off across all coverage levels, remaining stable even at low coverage.
@@
For practitioners aiming to identify the best overall selective classification setup that considers both the base classifier and the selector, one approach is to compare performance using the AURC metric. On the 1K-class datasets, EVA paired with $\Delta$-MDS-RLog achieves an AURC of 3.86, outperforming the DFN CLIP model with $\Delta$-KNN-RLog at 5.01. Despite similar NAURC values, EVA’s higher Im-1K base accuracy ($84.33\%$ vs. DFN CLIP’s $80.39\%$) makes it the preferred choice when considering both components. Intuitively, the optimal setup involves pairing the best selector (here, $\Delta$-MDS-RLog) with the most accurate base classifier.
For EVA, the ranking is reversed: $\Delta$-MDS-RLog achieves the best overall performance, followed by $\Delta$-KNN-RLog. This supports our hypothesis that MDS-based methods are particularly effective for supervised models due to the close connection between softmax classifiers and Gaussian Discriminant Analysis (Lee et al., 2018), which justifies the Gaussian assumptions used in MDS. In contrast, CLIP models trained with contrastive learning (Radford et al., 2021) do not satisfy these assumptions, making the nonparametric $\Delta$-KNN combination more suitable. The bottom row of Fig. 3 of the appendix confirms that $\Delta$-MDS-RLog yields the best risk-coverage behavior for supervised learning across all coverage levels. Comparing average NAURC on the full 1K-class datasets, $\Delta$-MDS-RLog with EVA is the top performing selector score, with a slightly better score (0.157) than the best performer on CLIP (0.163).
To verify that the performance gains of our methods stem from actual algorithmic improvements, rather than large-scale pretraining of CLIP or EVA potentially mitigating distribution shifts, we also evaluate on ResNet50 trained solely on ImageNet-1K. The results in appendix Table 7 show that our methods perform the best, consistent with earlier findings.
Semantic Shift Experiments.
For completeness, we also report results for experiments on datasets that are semantic shifts to ImageNet-1K in Appendix Table 8. In agreement with the covariate shift experiments, our proposed methods achieve the best performance on this benchmark.
@@ 5.2 Language Experiments
Table 2 presents results on the Amazon Reviews dataset. Unlike the vision tasks, the best-performing method is $\Delta$-MDS-MSP, followed closely by $\Delta$-MDS-RLog and $\Delta$-KNN-MSP. Since LISA (Yao et al., 2022) uses a softmax classification objective, the superiority of MDS-based selectors supports our hypothesis about their suitability for supervised models. Interestingly, MSP outperforms RLog in this domain, resulting in better performance when combined with $\Delta$-MDS. This highlights another important practical insight that the best linear combination often involves pairing the top-performing standalone distance-based and logit-based score.
@@ 5.3 Ablations
Table 5: Ablation results using DFN-CLIP on ImageNet-1K where the fraction of labeled samples used in feature computation for our proposed methods are varied.
0.1% 1% 10% 50% 100%
Method A N A N A N A N A N
$\Delta$-MDS-RLog - - 10.5 0.638 4.19 0.202 4.14 0.198 4.13 0.197
$\Delta$-KNN-RLog 4.81 0.245 4.58 0.229 4.17 0.200 3.98 0.188 3.98 0.187
Design Choices.
Table 4 summarizes several ablation experiments on the design choices of our proposed methods. First, we justify averaging the top-$k$ nearest neighbor distances in $\Delta$-KNN rather than using the $k$-th distance alone. This modification yields measurable gains, where average AURC improves from 6.03 to 5.89 and NAURC improves from 0.233 to 0.225. We also investigate various combinations of selector scores. For CLIP, $\Delta$-KNN-RLog remains the best across all configurations, outperforming both double-distance combinations (e.g., $\Delta$-MDS-$\Delta$-KNN) and double-logit combinations (e.g., MSP-RLog). Notably, pairing $\Delta$-KNN with RLog significantly outperforms pairing it with MSP, further validating RLog’s role as a strong logit-based complement.
Sample Efficiency.
Although our methods require a one-time feature computation step, this cost is amortized over all future inference runs as the resulting features can be cached. Nevertheless, to evaluate performance in low-data or low-computation resource regimes, we conducted ablations limiting the amount of labeled samples used. The results in Table 5 show that both methods are surprisingly stable. $\Delta$-KNN is especially robust, maintaining strong performance with as little as 0.1% of labeled data. As expected, $\Delta$-MDS degrades at the 1% level due to the difficulty of estimating per-class statistics with so few samples, and is not applicable at 0.1% (roughly 1 image per class). Importantly, $\Delta$-KNN-RLog continues to outperform RLog at 1% and matches it at 0.1% (see Table 1), indicating that our method is still preferable whenever even a small amount of labeled data is accessible.
@@ 6 Conclusion
We presented a framework for designing selector functions for selective classification, grounded in the Neyman–Pearson lemma. This reveals that the optimal selection score is a monotonic transformation of a likelihood ratio, which unifies several existing methods. We proposed two novel distance-based scores and their linear combinations with logit-based baselines. Experiments across vision and language demonstrate that our methods achieve state-of-the-art performance across diverse settings.
@@
While our focus has been on classification, the Neyman–Pearson framework is general and broadly applicable to other predictive tasks. Exploring selective prediction in settings where uncertainty plays a critical role, such as semantic segmentation and time series forecasting, presents promising future directions. Additionally, extending these ideas to generative models such as LLMs is another exciting avenue for future work.
Acknowledgements
This research / project is supported by A*STAR under its National Robotics Programme (NRP) (Award M23NBK0053). The authors also acknowledge support from Google (Google South Asia and South-East Asia Award).
Ethics Statement
This work does not involve human subjects, and it relies solely on publicly available models and datasets with all attributions provided. Based on the scope of our methods and results, we do not identify ethical issues that require special attention, and we are not aware of immediate harmful applications arising from this research.
Reproducibility Statement
We include the theoretical and experimental details necessary to reproduce our findings. Proofs and supporting discussions for all theoretical claims appear in appendix Sec. B and Sec. C. For implementation clarity, we provide pseudocode for our methods in appendix Sec. D and report experimental setup and hyperparameters in appendix Sec. E. Full source code with instructions is provided in the supplementary material to enable reproduction of the reported results.
LLM Usage
We used large language models to assist in formulating and checking mathematical proofs and to improve grammar and writing clarity. All outputs from the LLMs were reviewed by the authors before inclusion, and the authors take full responsibility for the paper’s content and the accuracy of the presented results.
@@ Acknowledgements
Appendix A Description of Baselines
In this section we provide a brief description of each baseline considered in this work.
Maximum Softmax Probability (MSP) (Hendrycks and Gimpel, 2017).
Given an input $\bm{x}$}, let the classifier output logits be $\{l^{(k)}\}_{k=1}^{K}$ and corresponding softmax probabilities $p_{\theta}(y=k|{\bm{x}})=\mathrm{softmax}(l^{(k)})$. The MSP score is defined as
$s_{\mathrm{MSP}}({\bm{x}})=\max_{k\in\{1,\dots,K\}}p_{\theta}(y=k|{\bm{x}}).$
MSP is commonly used as an OOD/confidence score: larger values indicate the model places high probability mass on a single class, and thus the input is treated as more “in-distribution” or “confident”. For selective classification, we threshold this score (and subsequent scores) to form the selector $g_{s,\gamma}({\bm{x}})=\mathbbm{1}[s_{\mathrm{MSP}}({\bm{x}})>\gamma]$.
Maximum Logit (MaxLogit) (Hendrycks et al., 2019).
The MaxLogit score is defined as
$s_{\mathrm{MaxLogit}}({\bm{x}})\;=\;\max_{k\in\{1,\dots,K\}}l^{(k)}$
Intuitively, larger values indicate that at least one class is assigned a large unnormalized score, and thus the model is more confident in its prediction.
Energy (Liu et al., 2020).
The energy score is defined from logits as the Helmholtz free energy
$E({\bm{x}})\;=\;-T\log\sum_{k=1}^{K}\exp\left(l^{(k)}/T\right),$
where $T>0$ is a temperature parameter (we set $T=1$). For selective classification we use the negative energy (so that larger values indicate more in-distribution-like inputs):
$s_{\mathrm{Energy}}({\bm{x}})\;=\;\log\sum_{k=1}^{K}\exp\left(l^{(k)}\right).$
Mahalanobis Distance (MDS) (Lee et al., 2018).
Let ${\bm{z}}=\phi({\bm{x}})$ denote the feature representation of $\bm{x}$} (e.g., penultimate features of a deep network). Using training data, we estimate the empirical mean feature $\mu_{i}$ for each class $i\in\{1,\dots,K\}$ and a shared (tied) covariance matrix $\Sigma$ across classes. The MDS score is then
$s_{\mathrm{MDS}}({\bm{x}})=\max_{i\in\{1,\dots,K\}}-({\bm{z}}-\mu_{i})^{\top}\Sigma^{-1}({\bm{z}}-\mu_{i}),$
i.e., the negative squared Mahalanobis distance to the closest class centroid. Intuitively, $s_{\mathrm{MDS}}({\bm{x}})$ is large when $\bm{x}$} lies in a high-density region of some class under the fitted Gaussian discriminant model, and small when $\bm{x}$} is far from all class clusters.
$k$-Nearest Neighbors (KNN) (Sun et al., 2022).
Let $\{{\bm{z}}_{j}\}_{j=1}^{n}$ denote the set of training features (we normalize the features). Define $r_{k}({\bm{z}})$ as the Euclidean distance from $\bm{z}$} to its $k$-th nearest neighbor among $\{{\bm{z}}_{j}\}$. The KNN score is
$s_{\mathrm{KNN}}({\bm{x}})=-\,r_{k}({\bm{z}}),$
so that inputs lying in denser regions of the training feature manifold (smaller neighbor distances) receive higher scores.
Raw Logits (RLog) (Liang et al., 2024).
RLog uses the confidence margin between the top two logits as a scale-robust confidence score. Let $l^{(1)}\geq l^{(2)}\geq\cdots\geq l^{(k)}$ be the logits sorted in descending order for a given input $\bm{x}$}. We define
$s_{\mathrm{RLog}}({\bm{x}})\;=\;l^{(1)}-l^{(2)}$
Intuitively, $s_{\mathrm{RLog}}({\bm{x}})$ is large when there is a clear “winner” class, and small when the classifier is uncertain between competing labels. Liang et al. (2024) argue that using a logit-space margin yields a more robust score under classifier miscalibration and post-hoc logit transformations (e.g., temperature scaling), since it depends only on the relative separation between the top classes.
Softmax Information Retaining Combination (SIRC) (Xia and Bouganis, 2022).
Let $S_{1}({\bm{x}})$ be a primary softmax-derived confidence score, and let $S_{2}({\bm{x}})$ be an auxiliary feature-based score. We choose $S_{1}$ to be MSP and $S_{2}$ to be the $L_{1}$ norm, in line with Xia and Bouganis (2022). SIRC combines these via
$s_{\mathrm{SIRC}}({\bm{x}})\;=\;-\bigl(S_{1}^{\max}-S_{1}({\bm{x}})\bigr)\Bigl(1+\exp\bigl(-b\,[S_{2}({\bm{x}})-a]\bigr)\Bigr),$
where $S_{1}^{\max}$ is the maximum attainable value of $S_{1}$ (e.g., $S_{1}^{\max}=1$ for MSP), and $a,b$ control how strongly $S_{2}$ influences the score. We follow Xia and Bouganis (2022) and set $a$ and $b$ using in-distribution statistics of $S_{2}$ ($a=\mu_{S_{2}}-3\sigma_{S_{2}}$ and $b=1/\sigma_{S_{2}}$).
Appendix B Proofs